OpenStudy (kittiwitti1):

I don't understand the answer: If the two square regions in the figures have the respective areas indicated in square yards, how many yards of fencing are needed to enclose the two regions? (Drawings below) options: a,\[4\sqrt{130}\]b,\[20\sqrt{10}\]c,\[24\sqrt{5}\]d,\[100\]e,\[105\sqrt{5}\]

OpenStudy (kittiwitti1):



OpenStudy (kittiwitti1):

Answer was \[\text{c) }24\sqrt{5}\]

OpenStudy (anonymous):

find the perimeter both sqaures. then add the perimeters

OpenStudy (kittiwitti1):

So \[\left(1\sqrt{5}+5\sqrt{5}\right)\times 4=6\sqrt{5}\times 4=24\sqrt{5}\]yes?

OpenStudy (kittiwitti1):

@ganeshie8

ganeshie8 (ganeshie8):

looks good !

OpenStudy (kittiwitti1):

Thank you @ganeshie8 :)

ganeshie8 (ganeshie8):

yw :)